OpenStudy (anonymous):

The area of a rectangle is 13 1/4 square feet. The length of this rectangle is 8 5/6 feet. What is the width? Help plz Fan + Medal a} 1 1/4 b} 1 5/8 c} 2/3 d} 1 1/2

OpenStudy (anonymous):

@aum ?

OpenStudy (anonymous):

@ShadowLegendX ?

OpenStudy (aum):

Area of a rectangle = Length x Width Width = Area / Length = \(13\frac 14 \div 8\frac 56 \) What is \(13\frac 14\) in fractions? What is \(8\frac 56\) in fractions?

OpenStudy (anonymous):

Those in inproper fractions?

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

13 1/4= 52/ 4 and 8 5/6 = 48/6

OpenStudy (aum):

\[ 13\frac 14 = \frac{13\times4+1}{4} = \frac{52+1}{4} = \frac{53}{4} \\ 8\frac 56 = ? \]

OpenStudy (anonymous):

Sorry about tht @aum i had to go lol